# Revisiting CMSSM with Non-Universal Gaugino Masses under Current Constraints

### v1

@@ III Results and discussion
  • Since an expected $r$ of signal strength greater than 1 is considered excluded by the experiment, it can be observed from the left panel of Fig. 7 that the expected $r$ values in Class  A and Class  C are relatively large. This indicates that direct searches for supersymmetric particles impose strong constraints on these classes. In contrast, such constraints are weaker for Class  B and the region where Class  A and Class  B intersect.
  • Combining the left panel of Fig. 6 and the left panel of Fig. 7, an interesting observation can be made. The samples in Class  A and Class  C with $\tilde{\chi}^{0}_{1}\tilde{\chi}^{\pm}_{1}$ coannihilation exhibit small expected $r$ values of signal strength. This indicates that, although direct searches for supersymmetric particles impose strong constraints on wino-like $\tilde{\chi}^{0}_{1}$ samples, these constraints become significantly weaker for wino-like samples with $\tilde{\chi}^{0}_{1}\tilde{\chi}^{\pm}_{1}$ coannihilation.
  • From the middle and right panels of Fig. 7, it can be observed that direct searches for supersymmetric particles impose strong constraints primarily on two scenarios: (1) when $\tilde{\tau}_{1}$ is light and the masses of $\tilde{\tau}_{1}$ and $\tilde{\chi}_{1}^{0}$ are not degenerate, and (2) when the masses of $\tilde{\chi}_{1}^{0}$ and $\tilde{\chi}_{1}^{\pm}$ are degenerate. In other cases, these constraints are significantly weaker.
  • From the right panel of Fig. 7, it can be observed that hadron colliders, including the LHC, HL-LHC, and HE-LHC, have very limited constraints on the compressed mass interval of the samples. Even with increased energy and luminosity, it remains challenging to cover all of these samples. In contrast, lepton colliders, such as the ILC and CLIC, demonstrate significantly better coverage in this region. They can effectively probe almost all samples with masses less than half of their collision energy, even with relatively modest integral luminosities. Overall, when the HL-LHC achieves an integral luminosity of $3\,\mathrm{ab}^{-1}$ and the CLIC1500 reaches $2.5\,\mathrm{ab}^{-1}$, all surviving samples will be fully covered by these experiments.

### v2

@@ III Results and discussion
  • Since an expected $r$ of signal strength greater than 1 is considered excluded by the experiment, it can be observed from the left panel of Fig. 7 that, compared to the samples with $M_{1}/M_{2}\lesssim 1.5$, the samples with $M_{1}/M_{2}\gtrsim 1.5$ are more likely to be excluded by experiments. There are exceptions for the sample with $M_{3}>0$ and $M_{1}/M_{2}\approx 1.5$ or the samples with $M_{3}<0$ and $M_{1}\approx 1.5M_{2}+200$ GeV. These samples are also hard to be excluded by current SUSY particles direct search experiment.
  • By combining the left panel of Fig. 5 with the left and middle panels of Fig. 7, an interesting observation can be made. Samples that are more easily excluded by current SUSY particles search experiments correspond to $\tilde{\chi}^{\pm}_{1}\tilde{\chi}^{0}_{1}$ coannihilation. Conversely, when the annihilation mechanism of the sample is $\tilde{\tau}\tilde{\chi}^{0}_{1}$ coannihilation, $\tilde{\chi}^{0}_{1}\tilde{\chi}^{\pm}_{1}$ coannihilation, or $\tilde{\chi}^{0}_{1}$ annihilation, the experimental constraints are much weaker. These observations indicate that the direct searches for SUSY particles experiment impose strong constraints on samples with non-degenerate mass of $\tilde{\tau}$ and $\tilde{\chi}^{0}_{1}$ ($m_{\tilde{\tau}_{1}}/m_{\tilde{\chi}^{0}_{1}}\gtrsim 1.1$) and degenerate mass of $\tilde{\chi}^{\pm}_{1}$ and $\tilde{\chi}^{0}_{1}$ ($m_{\tilde{\chi}^{\pm}_{1}}/m_{\tilde{\chi}^{0}_{1}}\approx 1$).
  • From the right panel of Fig. 7, it can be observed that hadron colliders, including the LHC, HL-LHC, and HE-LHC, have very limited constraints on the compressed mass interval of the samples. Even with increased energy and luminosity, it remains challenging to cover all of these samples. In contrast, lepton colliders, such as the ILC and CLIC, demonstrate significantly better coverage in this region. They can effectively probe almost all samples with masses less than half of their collision energy, even with relatively modest integral luminosities. In general, when the HL-LHC achieves an integral luminosity of $3\,\mathrm{ab}^{-1}$ and the CLIC1500 reaches $2.5\,\mathrm{ab}^{-1}$, all surviving samples will be fully covered by these experiments.
In this paper, we qualitatively discuss the sensitivity of future colliders, such as HL-LHC, HE-LHC, ILC, and CLIC, in probing the surviving parameter space based on experimental exclusion limits. In contrast, Ref.[72] uses Monte Carlo methods for a quantitative assessment of the HL-LHC and HE-LHC reach in different search channels, while Ref.[73] investigates the survival of benchmark points under current SUSY search constraints. Furthermore, the phenomenology of $(g-2)_{\mu}$, dark matter, and $B$ physics for specific benchmark points, such as $M_{1}=M_{2}$ and $M_{1}=5M_{2}$, has been discussed in detail in Ref. [74]. In contrast, our work extends this analysis by exploring scenarios with arbitrary ratios between $M_{1}$ and $M_{2}$. Additionally, we conduct a detailed study of dark matter annihilation mechanisms and their implications for direct detection and collider experiments. Compared to our previous work in Ref. [75], this study employs more sophisticated computational tools for parameter space analysis and includes a thorough discussion of the ability to cover the parameter space of future collider experiments.